# Advancements in Constitutive Model Calibration: Leveraging the Power of Full-Field DIC Measurements and In-Situ Load Path Selection for Reliable Parameter Inference

### v1

@@ 8 CRediT Authorship Contribution Statement
Appendix A Cruciform Specimen
The dimensions of the cruciform geometry simulated in this work are detailed in Fig. 1.
Figure 1: Abbreviated drawing of the cruciform specimen geometry. Dimensions are in inches.
Appendix B Retained Principal Components of the Nodal Displacements
The retained principal components at Node 1 in the load path tree (Fig. 2) are shown for both directional components X and Y in Figs. 1 and 2, respectively.
@@ Appendix B Retained Principal Components of the Nodal Displacements
Figure 1: The 5 retained principal components for the $X$ directional component of displacement after one load step along Axis A - i.e. Node 1 in the load path tree in Fig. 2
@@
Figure 2: The 5 retained principal components for the $Y$ directional component of displacement after one load step along Axis A - i.e. Node 1 in the load path tree in Fig. 2.
Appendix C EIG Approximation
\@fb@secFB
C.1 Challenges of Calculating the EIG
In each step of the ICC framework, a single load step is taken, data is collected for calibration, Bayesian inference is performed to obtain an approximation to the posterior distribution describing parameter uncertainty, and then the next load step is selected by calculating the expected information gain (EIG). The EIG is approximated via a nested Monte Carlo estimator for each candidate design $\xi$,
$\widehat{EIG}_{MC}(\xi)=\frac{1}{N}\sum_{i=1}^{N}\log\frac{f\left(\boldsymbol{y}_{i}\>|\>\boldsymbol{\theta}_{i,0},\xi\right)}{\frac{1}{M}\sum_{j=1}^{M}f\left(\boldsymbol{y}_{i}\>|\>\boldsymbol{\theta}_{i,j},\xi\right)},\quad\boldsymbol{\theta}_{i,j}\sim\pi(\boldsymbol{\theta}),\quad\boldsymbol{y}_{i}\sim f(\boldsymbol{y}\>|\>\boldsymbol{\theta}_{i,0},\xi),$ (16)
where the denominator is a Monte Carlo approximation of the evidence,
@@ C.1 Challenges of Calculating the EIG
The candidate design which yields the maximum EIG is expected to result in the most informative data for reducing parameter uncertainty.
Numerical underflow may occur when calculating the EIG as a result of one or more factors: 1) If the prior distribution is not sufficiently centered on the posterior of $\boldsymbol{\theta}_{i}$, this may lead to calculations for the evidence term (the inner loop of the EIG MC estimator) that are below machine precision [64], and 2) in the presence of high-dimensional data where the covariance-scaled squared difference of the measurements and model output (the exponential term in Eqn. (13)) are less than 1, the likelihood calculations (used to calculate the evidence) may be below machine precision. Both sources of underflow were seen in this work. The former was handled through importance sampling (details in C.2). The latter case is observed if the displacement field is used for calibration without dimension reduction; this was successfully resolved with the PCA dimension reduction process discussed in Sec. 3.2. In combination, these two approaches were successful in resolving issues with underflow.
\@fb@secFB
C.2 Importance Sampling
In the synthetic cruciform example, underflow was observed at the first adaptively selected load step (during the EIG calculation for load step two) when the sampling distribution in the EIG ($\pi(\boldsymbol{\theta})$) was not yet sufficiently centered on the posterior. As more data became available at later load steps, underflow was not observed. Furthermore, the occurrence of underflow was rare. Calculating the EIG with $M=1\times 10^{3}$ and $N=1\times 10^{4}$ resulted in three outer loop samples (.03% of the N outer loop samples) experiencing underflow for the EIG approxiamiton of an A load step, and 0 occurrences of underflow for the EIG approximation of a B load step. For a greater computational cost, increasing $M$ may aid in reducing the occurrence of underflow as more inner loop samples provide a better approximation of the evidence.
To overcome underflow when it occurred, importance sampling was used, whereby a sampling distribution alternative to the prior was chosen which contained more mass over the posterior of $\boldsymbol{\theta}_{i}$. Selection of a proper distribution to sample from is not a trivial task. Several different approaches have been proposed to calculate a sampling distribution for the EIG inner loop approximation of the evidence, two of which were studied for this work: 1) The Laplace-based importance sampling approach proposed by Beck et al. [64], and 2) the self-normalizing importance sampling approach proposed by Feng [65]. In addition to these, the analytical approach presented by Long et al. [66] was explored, which uses the Laplace approximation to estimate the evidence term of the inner loop.
It was found that the self-normalized distribution [65] was also susceptible to underflow for this problem, so this approach was not adopted. Using the Laplace-based biasing distribution [64] for all $N$ outer loop iterations would be prohibitively expensive (a single iteration of the outer loop requires 8 seconds as opposed to 8 milliseconds when sampling from the prior). Likewise, an analytical calculation of the inner loop [66] was computationally expensive. Thus, in this work, a hybrid sampling approach was used: The prior distribution was sampled from by default, but if underflow occurred, the Laplace-based importance sampling approach [64] was reverted to in order to overcome underflow for that iteration of the outer loop. Since underflow occurred for such a dearth of outer loop iterations, the hybrid approach was feasible and successfully overcame the issue with underflow with an unremarkable increase in computational time.
Appendix D Linear Transformation of a Random Variable
For some random variable $Q$ that has expected value $\mathbb{E}[Q]$ and variance $\text{Var}[Q]$, the transformation of $Q$ via $AQ$, where $A$ is a constant matrix $A\in\mathbb{R}^{i,j}$ results in  [67]
$\mathbb{E}[AQ]=A\mathbb{E}[Q],$ (18)
$\text{Var}[AQ]=A\text{Var}[Q]A^{T}.$ (19)
Nodal displacement data along with global load values were used to calibrate the material model parameters through Bayesian inference. The data $\boldsymbol{y}=[\text{disp}_{i},\text{load}_{i}]\,i\in[X,Y]$ are considered as random variables in this context. Thus, when the displacement data is dimensionally reduced—through PCA via the process discussed in Sec. 3.2—a linear transformation takes places such that $\text{disp}_{i}^{\text{PCA}}=\text{disp}_{i}V_{i}^{*}$ for $i\in[X,Y]$, where $V_{i}^{*}$ is the PCA basis of the retained principal components for the given directional component. After the transformation, $\boldsymbol{y}=[\text{disp}_{i}^{\text{PCA}},\text{load}_{i}],\,i\in[X,Y]$.
\@fb@secFB
D.1 Likelihood Construction
The likelihood for the data is modeled as a product of the likelihood of each of the 4 data sets contained in $\boldsymbol{y}$ (load and displacement for $X$ and $Y$ directional components) under the assumption of independence,
$f\left(\boldsymbol{y}\>|\>\boldsymbol{\theta},\boldsymbol{\xi}\right)=\prod_{h}\prod_{i}f\left(\boldsymbol{y}_{h,i}\>|\>\boldsymbol{\theta},\boldsymbol{\xi}\right),\quad h\in[\text{disp}^{\text{PCA}},\text{load}],\quad i\in[X,Y].$ (20)
Each data set is modeled with a multivariate normal distribution with i.i.d. Gaussian error. The multivariate normal distribution for each data set is characterized by its mean $\mathbb{E}[\boldsymbol{y}_{h,i}]=\tilde{g}(\boldsymbol{\theta})_{h,i}$ and variance $\text{Var}[\boldsymbol{y}_{h,i}]=\Psi_{h,i}$ such that $\boldsymbol{y}_{h,i}\sim\mathcal{N}\left(\mathbb{E}[\boldsymbol{y}_{h,i}],\text{Var}[\boldsymbol{y}_{h,i}]\right)=\mathcal{N}\left(\tilde{g}(\boldsymbol{\theta})_{h,i},\Psi_{h,i}\right)$. Thus, the likelihood for each independent data set is,
$f\left(\boldsymbol{y}_{h,i}\>|\>\boldsymbol{\theta},\boldsymbol{\xi}\right)$ $=\mathcal{N}_{N_{h,i}}\left(\boldsymbol{y}_{h,i};\tilde{g}(\boldsymbol{\theta})_{h,i},\Psi_{h,i}\right),$ (21)
$h\in[\text{disp}^{\text{PCA}},\text{load}],\,i\in[X,Y].$
The PCA-transformed displacement field is obtained via a linear transformation, $\text{disp}^{PCA}_{i}=\text{disp}_{i}V_{i}^{*}$, and is distributed as $\text{disp}^{PCA}_{i}\sim\mathcal{N}_{N_{\text{disp}^{PCA},i}}\left(\tilde{g}(\boldsymbol{\theta})_{\text{disp}^{\text{PCA}},i},\,V^{*T}_{i}\text{Var}[\text{disp}_{i}]V_{i}^{*}\right)$ for $i\in[X,Y]$. Recall, the surrogate models $\tilde{g}(\boldsymbol{\theta})$ take a parameter vector as input an return PCA singular values for the displacement field and load values. The noise variance for the displacement is defined in the field space, so it must be transformed according to Eqn. (19) in order to account for the PCA transformation of the displacement field.
Appendix E Parameter Correlation
See Tables 1 and 2 for posterior correlation of the AAAAA and ABABA load paths, respectively.
@@ Appendix E Parameter Correlation
Table 1: Element-wise posterior correlation for load path AAAAA.
@@
$A$ 0.50 1.
$n$0.63 -0.97 1.
$a$0.17 0.14 -0.02 1.
Table 2: Element-wise posterior correlation for load path ABABA.
Appendix F Additional Figures for Load Path AAAAA Calibration
Additional figures for the load path AAAAA calibration are contained here. These include the synthetic displacement data in F.1, the posterior probability over the PCA singular values in F.2, the posterior predictive distribution over the displacement field for the ABBBA load path in F.3 and a table reporting the log predictive density for various load paths in F.4.
@@ Appendix F Additional Figures for Load Path AAAAA Calibration
F.1 Synthetic Nodal Displacement Data
See Figs. 1 and 2 for the synthetic nodal displacement data used for calibration with an AAAAA load path.
@@ F.1 Synthetic Nodal Displacement Data
Figure 1: Synthetic DIC data for AAAAA load path. The X component of displacement is shown in microns for load steps 1-5 in plots (a)-(e), respectively.
@@
Figure 2: Synthetic DIC data for AAAAA load path. The Y component of displacement is shown in microns for load steps 1-5 in plots (a)-(e), respectively.
@@
F.2 Posterior probability of PCA Singular Values
See Fig. 3 for a visualization of the posterior probability over the PCA singular values for the AAAAA load path.
Figure 3: Posterior probability over the 5 retained PCA singular values for the AAAAA load path after load step 5 for both the $X$ (top row) and $Y$ (bottom row) directional components. The singular values at the MAP parameter estimate is shown with a dotted black line, and the 95% CI of the distributions are shown with dashed black lines, indicating the interval in which the singular values have a 95% probability of residing within.
\@fb@secFB
F.3 Validation with ABBBA Load Path
Validation of the model calibrated with an AAAAA load path was performed by making predictions about a new load path ABBBA. Figs. 15 (X directional component) and 4 (Y directional component) show samples from the posterior predictive distribution Eqn. (15) over the displacement field. The left-most column shows the observed ABBBA displacement data and the next three columns show three different draws from the posterior predictive distribution.
In Fig. 5, the distribution of the field is shown over Line 2 from Fig. 16 and is represented through the 95% CI (dashed lines). The distribution is shown for all 5 load steps (LS). The left plots show the 95% CI along with the observed data, and the right plots show the difference between the 95% CI and the observed data. The solid line represents the observations (centered at 0 difference) and the difference between the observations and the CIs are shown with dashed lines. Everywhere the observations fall within the 95% CI coincides with locations where the upper bound difference is positive and the lower bound difference is negative; these regions are colored orange.
Figure 4: Samples from the posterior predictive distribution of an AAAAA load path is shown for the $Y$ directional component of the displacement field for a new ABBBA load path. The left column shows the observed data for each of the 5 load steps (LS) and the three right columns show three different predictive samples.
Figure 5: The posterior predictive distribution (from an AAAAA load path) over Line 2 displacements (Fig. 16) is shown for new load path ABBBA. The two left columns show the 95% CI of the distribution as well as the observations for the $X$ and $Y$ directional components for all 5 load steps (LS). The two right columns show the difference between the 95% CI and the observations. Regions where the observations fall outside the 95% CI are colored orange, and 0 is marked with a red line. The x-axis is the normalized location along Line 2.
\@fb@secFB
F.4 Log Predictive Density
An additional comparison can be made between the model calibration with an AAAAA load path vs an ABABA load path by comparing their predictive accuracy given a point estimate of the fitted model, $\hat{\boldsymbol{\theta}}$—namely, the MAP parameter estimate which yields the best-fit log predictive density (a.k.a. log-likelihood) [68], $\log\pi(\boldsymbol{y}^{pred}\>|\>\boldsymbol{\theta}^{\text{MAP}})$. This predictive score summarizes the fit of the model to the data and is proportional to the mean squared error in this case since a normal model and constant variance are assumed. Table 1 reports the log predictive density of the AAAAA- and ABABA-calibrated models for 16 new data sets that were not seen in calibration for the indicated load paths. For all but a few cases, the AAAAA-calibrated model yielded best-fit predictive densities that were greater than the ABABA-calibrated model.
Prediction Path Calibration Path Best-fit Log Predictive Density
$\boldsymbol{(1\times 10^{3})}$
[…]
ABABA 113,268
AAAAB AAAAA 113,907
ABABA 113,882
AAABA AAAAA 114,291
ABABA 114,257
AAABB AAAAA 114,115
ABABA 114,116
AABAA AAAAA 114,206
ABABA 114,184
AABAB AAAAA 114,346
ABABA 114,340
AABBA AAAAA 114,476
ABABA 114,463
AABBB AAAAA 114,169
ABABA 114,116
ABAAA AAAAA 113,854
ABABA 113,752
ABAAB AAAAA 114,331
ABABA 114,344
ABABA AAAAA 114,515
ABABA 114,648
ABABB AAAAA 114,293
ABABA 114,274
ABBAA AAAAA 114,518
ABABA 114,492
ABBAB AAAAA 114,173
ABABA 114,168
ABBBA AAAAA 114,153
ABABA 114,150
ABBBB AAAAA 113,997
ABABA 113,858
Table 1: Best-fit log predictive density.
Appendix G Additional Figures for Load Path ABABA Calibration
Additional figures for the load path ABABA calibration are contained here. These include the synthetic displacement data in G.1, the posterior probability in G.2 and the posterior predictive probability over the new load path ABBBA in G.3.
\@fb@secFB
G.1 Synthetic Nodal Displacement Data
See Figs. 1 and 2 for the synthetic nodal displacement data used for calibration with an ABABA load path.
@@ G.1 Synthetic Nodal Displacement Data
Figure 1: Synthetic DIC data for ABABA load path. The X component of displacement is shown in microns for load steps 1-5 in plots (a)-(e), respectively.
@@
Figure 2: Synthetic DIC data for ABABA load path. The Y component of displacement is shown in microns for load steps 1-5 in plots (a)-(e), respectively.
\@fb@secFB
G.2 Posterior Probability
The posterior probability contours for the ABABA load path can be found in Fig. 3 as well the posterior probability over the load in Fig. 4, PCA singular values in Fig. 5, and the displacement field in Figs. 6 & 7.
Figure 3: The posterior contour for load path ABABA. Marginal posterior densities are plotted on the diagonals and joint-marginal densities on the off-diagonals. $\boldsymbol{\theta}^{true}$ is indicated with a dashed pink line and star and the 95% CI with dashed black lines.
Figure 4: Posterior probability over the load for the ABABA load path for both the $X$ (left column) and $Y$ (right column) directional components. The load at the MAP parameter estimate is shown with a dotted black line, and the 95% CI of the distribution over the load is shown with dashed black lines, indicating the interval in which the load values have a 95% probability of residing within.
Figure 5: Posterior probability over the 5 retained PCA singular values for the ABABA load path after load step 5 for both the $X$ (top row) and $Y$ (bottom row) directional components. The singular values at the MAP parameter estimate is shown with a dotted black line, and the 95% CI of the distributions are shown with dashed black lines, indicating the interval in which the singular values have a 95% probability of residing within.
@@ G.2 Posterior Probability
Figure 6: The observed displacement field generated with $\boldsymbol{\theta}^{true}$ and added noise for the (a) X directional component and (b) Y directional component for the ABABA load path at the final load step reported in microns. Lines (1) and (2) indicate the space over which the posterior probability of the displacement field is plotted in Fig. 7.
@@
Figure 7: Posterior probability over the nodal displacements for the ABABA load path across the two lines indicated in Fig. 6. The $X$ (left column) and $Y$ (right column) directional components are shown along with Line 1 (top row) and Line 2 (bottom row). The 95% CI of the distribution over the load is shown with dashed black lines. Given the relative scale of the displacements and the low uncertainty, the 95% CI interval appears as a single line. The x-axis is the normalized location across each line scan.
@@
G.3 Posterior Predictive Probability
Validation of the model calibrated with an ABABA load path was performed by making predictions about a new load path ABBBA. The posterior predictive distribution—Eqn. (15)—is shown over the load in Fig. 8 and over the displacement field in Figs. 9-12.
@@ G.3 Posterior Predictive Probability
Figure 8: The posterior predictive distribution of the ABABA load path is shown for the new ABBBA load path over the global load values for the $X$ (left column) and $Y$ (right column) directional components at each load step. The observed data is shown with a solid red line, the load at the MAP parameter estimate is shown with a dotted black line, and the 95% highest density CI is shown with dashed black lines.
@@
Figure 9: Samples from the posterior predictive distribution of the ABABA load path is shown for the $X$ directional component of the displacement field for a new ABBBA load path. The left column shows the observed data for each of the 5 load steps (LS) and the three right columns show three different predictive samples.
@@
Figure 10: Samples from the posterior predictive distribution of the ABABA load path is shown for the $Y$ directional component of the displacement field for a new ABBBA load path. The left column shows the observed data for each of the 5 load steps (LS) and the three right columns show three different predictive samples.
@@
Figure 11: The posterior predictive distribution from the ABABA load path is displayed over the Line 1 displacements (Fig. 16) for a new ABBBA load path. The two left columns show the 95% CI of the distribution as well as the observations for the $X$ and $Y$ directional components for all 5 load steps (LS). The two right columns show the difference between the 95% CI and the observations. Regions where the observations fall outside the 95% CI are colored orange, and 0 is marked with a red line. the x-axis is the normalized location along Line 1.

### v3

@@ 8 CRediT Authorship Contribution Statement
Appendix A Surrogate Parity Plots
Parity plots are a graphical tool used here to assess the performance of the surrogate models by comparing predicted values against observed values. Here, there observed data is FE model output that was reserved for testing. In these plots, the x-axis represents the observed values, while the y-axis displays the predicted values. Ideally, if the model’s predictions are accurate, the points will lie along a 45-degree line, known as the line of perfect agreement. Deviations from this line indicate discrepancies between the predicted and observed values. The parity plots shown in this section assess the performance of the surrogate models at the final load step of an ABABA load path for the $X$ directional component. Fig. 1 shows parity plots for the displacement PCA singular values (SV), and Fig. 2 contains parity plots for the global load values, which are the two quantities the surrogate models predict. Fig. 2 also contains a parity plot for the displacement field, which was obtained by reconstructing field data from the surrogate prediction of the PCA singular values; any error between the true values and predicted values in this plot is due to the surrogate error as well as error originating from the PCA reconstruction. In all cases, there is very good alignment between the predicted values and true values, and there is not evidence of systematic discrepancy between the two.
@@ Appendix A Surrogate Parity Plots
Figure 1: Parity plots for the predicted PCA singular values (SV) from the surrogate model at the final load step of an ABABA load path for the $X$ directional component. The observed data is plotted on the x-axis and the surrogate prediction is plotted on the y-axis. The red 45-degree line is the line of perfect agreement.
[…]
Figure 2: Parity plots for the surrogate-predicted global load values (a) and reconstructed displacement fields (b) at the final load step of an ABABA load path for the $X$ directional component. The observed data is plotted on the x-axis and the surrogate prediction is plotted on the y-axis. The red 45-degree line is the line of perfect agreement.
Appendix B Posterior Contours
Fig. 1 shows the posterior contours for the AAAAA and ABABA calibrations obtained via the Laplace approximation (top row) as well as through MCMC simulation (bottom row) as a validation. MCMC was performed with an adaptive Metropolis-Hastings sampling algorithm; 120,000 samples were obtained with 20,000 samples discarded as burn-in. The chain was thinned by a factor of 10 to produce independent samples. The marginal distributions are shown on the diagonals plots along with the 95% credible interval (CI) (dashed black lines) and $\boldsymbol{\theta}^{true}$ (pink dashed line). The off-diagonals show the joint marginal posterior distributions along with $\boldsymbol{\theta}^{true}$ indicated by a pink star. Additional quantitative validation is contained in Tables 1 - 3. Table 1, reports posterior summaries obtained via the Laplace approximation as well as MCMC. Tables 2 and 3 report posterior correlations obtained via the Laplace approximation and MCMC, respectively, for both the AAAAA and ABABA calibrations. The posterior correlation describes the trade-off relationship between different parameters. Values closer to 1 (perfect correlation) indicate a stronger relationship between parameters than those closer to 0 (independence). The reported values correspond to the posterior contours shown in Fig. 1, where elongated contours suggest a correlation between the given parameters, while circular contours suggest little to no parameter correlation.
[…]
[figure]style=plain,subcapbesideposition=top,font=small
@@ Appendix B Posterior Contours
Figure 1: The posterior contour from the (a) AAAAA and (b) ABABA calibrations obtained via the Laplace approximation are shown on the top row. The bottom row shows posterior contours obtained via MCMC as a validation for (c) AAAAA and (d) ABABA. Marginal posterior densities are plotted on the diagonals and joint-marginal densities on the off-diagonals. $\boldsymbol{\theta}^{true}$ is indicated with a dashed pink line and star and the 95% CI with dashed black lines.
Load path Method $\boldsymbol{\mathbb{E}_{\boldsymbol{\theta}\>|\>\boldsymbol{y},\boldsymbol{\xi}}}$ $\boldsymbol{\mathbb{V}_{\boldsymbol{\theta}\>|\>\boldsymbol{y},\boldsymbol{\xi}}}$
$\boldsymbol{(1\times 10^{-4})}$
@@
MCMC [42.50, 13.65, 14.29, 11.19] [ 1.6, 2.5, 15.9, 5.2 ]
ABABA Laplace [42.53, 13.60, 14.31, 11.17] [ 2.2, 193.6, 580.2, 9.6 ]
MCMC [42.53, 13.62, 14.29, 11.17] [ 2.2, 194.7, 574.0, 9.6 ]
Table 1: Posterior summaries obtained via a Laplace approximation and MCMC simulation using data collected from the respective load paths.
@@
$A$ 0.52 1.
$n$0.63 -0.98 1.
$a$0.19 0.10 0.01 1.
Table 2: Element-wise posterior correlation for load path AAAAA (left) and ABABA (right) obtained via the Laplace approximation.
$\sigma_{y}$ $A$ $n$ $a$
[…]
$A$ 0.35 1.
$n$0.56 -0.64 1.
$a$0.18 0.56 0.20 1.
$\sigma_{y}$ $A$ $n$ $a$
[…]
$A$ 0.52 1.
$n$0.63 -0.98 1.
$a$0.18 0.10 0.01 1.
Table 3: Element-wise posterior correlation for load path AAAAA (left) and ABABA (right) obtained with MCMC.
@@
B.1 Displacement Posterior Predictive Fields
Figure 15 in Sec. 4.3 displays draws from the posterior predictive distribution for the displacement field to establish that the observed data is reasonable under the calibrated model. Here, complementary information is shown, which is the difference between the observed field and the predictive fields. From this, the error at each load step and for each predictive field can be observed. There are no observed systematic discrepancies and the error is similar among the multiple draws.
Figure 2: The error for three samples from the AAAAA posterior predictive distribution are shown for the $X$ directional component of the displacement field for the new ABBBA load path. The left column shows the observed data for each of the 5 load steps (LS) and the three right columns shows the error between three different predictive samples and the validation data.
Appendix C Log Predictive Density
An additional comparison can be made between the model calibration with an AAAAA load path vs an ABABA load path by comparing their predictive accuracy given a point estimate of the fitted model, $\hat{\boldsymbol{\theta}}$—namely, the MAP parameter estimate which yields the best-fit log predictive density (BFLPD) (a.k.a. log-likelihood), $\log\pi(\boldsymbol{y}^{pred}\>|\>\boldsymbol{\theta}^{\text{MAP}})$ [75]. This predictive score summarizes the fit of the model to the data and is proportional to the mean squared error in this case since a normal model and constant variance are assumed. Table 4 reports the log predictive density of the AAAAA- and ABABA-calibrated models for 16 new data sets that were not seen in calibration for the indicated load paths. For all but a few cases, the AAAAA-calibrated model yielded best-fit predictive densities that were greater than the ABABA-calibrated model.
Prediction Path Calibration Path BFLPD
$\boldsymbol{(1\times 10^{3})}$
[…]
ABABA 114,072
AAAAB AAAAA 113,911
ABABA 113,917
AAABA AAAAA 114,304
ABABA 114,287
AAABB AAAAA 114,114
ABABA 114,109
AABAA AAAAA 114,206
ABABA 114,206
AABAB AAAAA 114,345
ABABA 114,345
AABBA AAAAA 114,491
ABABA 114,489
AABBB AAAAA 114,180
ABABA 114,177
ABAAA AAAAA 113,856
ABABA 113,854
ABAAB AAAAA 114,343
ABABA 114,338
ABABA AAAAA 114,510
ABABA 114,270
ABABB AAAAA 114,300
ABABA 114,296
ABBAA AAAAA 114,520
ABABA 114,519
ABBAB AAAAA 114,183
ABABA 114,179
ABBBA AAAAA 114,167
ABABA 114,162
ABBBB AAAAA 114,022
ABABA 114,000
Table 4: Best-fit log predictive density.
Supplemental Material: Additional Figures and Tables
Appendix Supp 1 Cruciform Specimen
The dimensions of the cruciform geometry simulated in this work are detailed in Fig. S1.
Figure S1: Abbreviated drawing of the cruciform specimen geometry. Dimensions are in inches.
Appendix Supp 2 Retained Principal Components of the Nodal Displacements
The retained principal components at Node 1 in the load path tree (Fig. 2 in the main body of the text) are shown for both directional components X and Y in Figs. S2 and S3, respectively.
Node 1 Principal Components for X Directional Component
@@ Appendix C Log Predictive Density
Figure S2: The 5 retained principal components for the $X$ directional component of displacement after one load step along the X (horizontal) Axis (load step A) - i.e. Node 1 in the load path tree in Fig. 2
Node 1 Principal Components for Y Directional Component
@@
Figure S3: The 5 retained principal components for the $Y$ directional component of displacement after one load step along the X (horizontal) Axis (load step A) - i.e. Node 1 in the load path tree in Fig. 2.
Appendix Supp 3 Additional Figures and Tables from the Calibrations
Additional figures and tables from the calibrations are contained here. These include the synthetic displacement data in Supp 3.1, the posterior probability over the QoIs in Supp 3.2 and the posterior predictive distribution over the QoIs in Supp 3.3.
@@
Supp 3.1 Synthetic Nodal Displacement Data
AAAAA Data X Directional Component
@@
Figure S4: Synthetic DIC data for the AAAAA load path. The X component of displacement is shown in microns for load steps 1-5 in plots (a)-(e), respectively.
AAAAA Data Y Directional Component
@@
Figure S5: Synthetic DIC data for the AAAAA load path. The Y component of displacement is shown in microns for load steps 1-5 in plots (a)-(e), respectively.
ABABA Data X Directional Component
@@
[…]
Figure S6: Synthetic DIC data for the ABABA load path. The X component of displacement is shown in microns for load steps 1-5 in plots (a)-(e), respectively.
ABABA Data Y Directional Component
[…]
Figure S7: Synthetic DIC data for the ABABA load path. The Y component of displacement is shown in microns for load steps 1-5 in plots (a)-(e), respectively.
\@fb@secFB
Supp 3.2 Posterior Probability Over the QoIs
The posterior probability is visualized over the QoIs (load and nodal displacements) for both the AAAAA and ABABA calibrations in this section.
AAAAA Posterior Over Load
ABABA Posterior Over Load
Figure S8: Posterior probability over the load for the AAAAA (top row) and ABABA (bottom row) calibrations are shown for both the $X$ (left column) and $Y$ (right column) directional components. The load at the MAP parameter estimate is shown with a dotted black line, and the 95% CI of the distribution over the load is shown with dashed black lines, indicating the interval in which the load values have a 95% probability of residing within.
AAAAA Load Step 5 Data
[…]
Figure S9: The observed displacement field at the final load step of load path AAAAA generated with $\boldsymbol{\theta}^{true}$ and added noise for the (a) X directional component and (b) Y directional component reported in microns. Lines (1) and (2) indicate the space over which the posterior probability of the displacement field is plotted in Fig. S10.
AAAAA Posterior Over Displacement: Lines 1 and 2
@@
Figure S10: Posterior probability over the nodal displacements for the AAAAA calibration across the two lines indicated in Fig. S9. The $X$ (left column) and $Y$ (right column) directional components are shown along with Line 1 (top row) and Line 2 (bottom row). The 95% CI of the distribution over the load is shown with dashed black lines. Given the relative scale of the displacements and the low uncertainty, the 95% CI interval appears as a single line. The x-axis is the normalized location across each line scan.
ABABA Load Step 5 Data
[…]
Figure S11: The observed displacement field at the final load step of load path ABABA generated with $\boldsymbol{\theta}^{true}$ and added noise for the (a) X directional component and (b) Y directional component reported in microns. Lines (1) and (2) indicate the space over which the posterior probability of the displacement field is plotted in Fig. S12.
ABABA Posterior Over Displacement: Lines 1 and 2
[…]
[]Refer to caption
[…]
Figure S12: Posterior probability over the nodal displacements for the ABABA calibration across the two lines indicated in Fig. S11. The $X$ (left column) and $Y$ (right column) directional components are shown along with Line 1 (top row) and Line 2 (bottom row). The 95% CI of the distribution over the load is shown with dashed black lines. Given the relative scale of the displacements and the low uncertainty, the 95% CI interval appears as a single line. The x-axis is the normalized location across each line scan.
@@
Supp 3.3 Validation with ABBBA Load Path
Validation of the calibrated models was performed by making predictions about a new load path ABBBA. Samples from the posterior predictive distribution—Eqn. (15) in the main body of the text—are shown over the load and displacement field for both the AAAAA and ABABA-calibrated models. The left-most column of Figs. S14, S15, S21 and S22 shows the observed displacement data for load path ABBBA and the next three columns show three different draws from the posterior predictive distribution. The distribution of the field over the line plots in Fig. S18 is shown in Figs. S19, S20, S25 and S26 for both the AAAAA and ABABA calibration. The left plots show the 95% PI along with the observed data, and the right plots show the difference between the 95% PI and the observed data. The solid line represents the observations (centered at 0 difference) and the difference between the observations and the PIs are shown with dashed lines. Everywhere the observations fall within the 95% PI coincides with locations where the upper bound difference is positive and the lower bound difference is negative; these regions are colored orange.
Posterior Predictive Distribution Over ABBBA Load for the (Left) AAAAA— and (Right) ABABA—Calibrated Model
@@
Figure S13: The posterior predictive distribution of the (left) AAAAA— and (right) ABABA—calibrated model is shown for the new ABBBA load path over the global load values for the $X$ and $Y$ directional components at each load step. The observed data is shown with a solid red line, the load at the MAP parameter estimate is shown with a dotted black line, and the 95% highest density PI is shown with dashed black lines.
AAAAA Posterior Predictive Distribution Over ABBBA X Displacement
@@
Figure S14: Samples from the posterior predictive distribution of the AAAAA—calibrated model is shown for the $X$ directional component of the displacement field for a new ABBBA load path. The left column shows the observed data for each of the 5 load steps (LS) and the three right columns show three different predictive samples.
AAAAA Posterior Predictive Distribution Over ABBBA Y Displacement
@@
Figure S15: Samples from the posterior predictive distribution of the AAAAA—calibrated model is shown for the $Y$ directional component of the displacement field for a new ABBBA load path. The left column shows the observed data for each of the 5 load steps (LS) and the three right columns show three different predictive samples.
Error of AAAAA Posterior Predictive Distribution Over ABBBA X Displacement
Figure S16: The error for three samples from the AAAAA posterior predictive distribution are shown for the $X$ directional component of the displacement field for the new ABBBA load path. The left column shows the observed data for each of the 5 load steps (LS) and the three right columns shows the error between three different predictive samples and the validation data.
Error of AAAAA Posterior Predictive Distribution Over ABBBA Y Displacement
Figure S17: Samples from the posterior predictive distribution of the AAAAA—calibrated model is shown for the $Y$ directional component of the displacement field for a new ABBBA load path. The left column shows the observed data for each of the 5 load steps (LS) and the three right columns show three different predictive samples.
ABBBA Load Step 5 Data
[…]
Figure S18: The observed displacement field at the final load step of load path ABBBA generated with $\boldsymbol{\theta}^{true}$ and added noise for the (a) X directional component and (b) Y directional component reported in microns. Lines (1) and (2) indicate the space over which the posterior probability of the displacement field is plotted in the following figures.
AAAAA Posterior Predictive Distribution Over ABBBA Displacements: Line 1
@@
Figure S19: The posterior predictive distribution of the AAAAA—calibrated model over Line 1 displacements (Fig. S18) is shown for new load path ABBBA. The two left columns show the 95% PI of the distribution as well as the observations for the $X$ and $Y$ directional components for all 5 load steps (LS). The two right columns show the difference between the 95% PI and the observations. Regions where the observations fall outside the 95% PI are colored orange, and 0 is marked with a red line. The x-axis is the normalized location along Line 1.
AAAAA Posterior Predictive Distribution Over ABBBA Displacements: Line 2